# A thermonuclear supernova interacting with hydrogen- and helium-deficient circumstellar material

### v1

@@
Identifying the progenitors of thermonuclear supernovae (Type Ia supernovae; SNe Ia) remains a key objective in contemporary astronomy. The rare subclass of SNe Ia that interacts with circumstellar material (Type Ia-CSM) allows for studies of the progenitor’s environment before explosion, and generally favours single-degenerate progenitor channels. The case of SN Ia-CSM PTF11kx clearly connected thermonuclear explosions with hydrogen-rich CSM-interacting events, and the more recent SN 2020eyj connected SNe Ia with helium-rich companion progenitors. Both of these objects displayed delayed CSM interaction which established their thermonuclear nature. Here we present a study of SN 2020aeuh, a Type Ia-CSM with delayed interaction. We analyse photometric and spectroscopic data that monitor the evolution of SN 2020aeuh and compare its properties with those of peculiar SNe Ia and core-collapse SNe. At early times, the evolution of SN 2020aeuh resembles a slightly overluminous SN Ia. Later, the interaction-dominated spectra develop the same pseudocontinuum seen in Type Ia-CSM PTF11kx and SN 2020eyj. However, the later-time spectra of SN 2020aeuh lack hydrogen and helium narrow lines. Instead, a few narrow lines could be attributed to carbon and oxygen. We fit the pseudobolometric light curve with a CSM-interaction model, yielding a CSM mass of M☉. We propose that SN 2020aeuh was a Type Ia supernova that eventually interacted with a dense medium which was deficient in both hydrogen and helium. Whereas previous SNe Ia-CSM constitute our best evidence for nondegenerate companion progenitors, the CSM around SN 2020aeuh is more difficult to understand. We include a hydrodynamical simulation for a double-degenerate system to showcase how the dynamical evolution of such a progenitor scenario could produce the CSM observed around SN 2020aeuh. It is clear that SN 2020aeuh challenges current models for stellar evolution leading up to a SN Ia explosion.
@@
Supernovae (SNe) interacting with circumstellar material (CSM) are classified based on their spectra, which are dominated by narrow ( km s-1) emission lines (e.g. Gal-Yam 2017). These emission lines are interpreted as signs of shock interaction between fast-moving SN ejecta and slow-moving CSM (see, e.g. Smith 2017, for a review). The most commonly observed subtypes of interacting SNe are Type IIn (Schlegel 1990; Nyholm et al. 2020) and Type Ibn (Hosseinzadeh et al. 2017), which have spectra dominated by hydrogen (H) and helium (He) emission lines, respectively. Recently, the classes of H/He-poor CSM-interacting stripped-envelope core-collapse supernovae (SNe Icn; Gal-Yam et al. 2022; Perley et al. 2022) and even a proposed SN Ien (Schulze et al. 2024), were also discovered.
@@ 2.1 Discovery, initial observations and classification
SN 2020aeuh (also known as ATLAS20bkbb, ZTF20acyroke) was first reported to the Transient Name Server (TNS111https://wis-tns.weizmann.ac.il) in January 2021 by the Asteroid Terrestrial-impact Last Alert System (ATLAS) collaboration (Tonry et al. 2021), with their first detection from Christmas day 2020 December 25 (UTC dates are used throughout this paper; = 59208.6 MJD) at mag in the cyan band.
@@ 2.2 Photometry
Following the discovery, we obtained regular follow-up photometry with P48 until SN 2020aeuh faded below the detection limit, 180 days after the first detection. Initial follow-up observations were performed in the and bands, with later phases also covered in the band. This was accomplished with the ZTF camera (Dekany et al. 2020) on the P48. The , , and band P48 data include points on 88, 86, and 23 separate nights, respectively.
The photometry from the P48 come from the ZTF forced-photometry service (Masci et al. 2019, 2023) and from the P60 come from the pipeline described in Fremling et al. (2016). The photometric magnitudes of SN 2020aeuh are listed in Table 1. Additionally, we make use of the forced-photometry service from the ATLAS survey (Tonry et al. 2018; Smith et al. 2020; Shingles et al. 2021), which contains valuable photometry in the and bands.
@@ 2.3 Spectroscopy
The public SN Ia classification spectrum of SN 2020aeuh mentioned above was obtained 24 days after discovery (Weil et al. 2021). The SN never quite reached apparent magnitude 19, which is the cutoff for the Bright Transient Survey (BTS; Fremling et al. 2020; Perley et al. 2020). It was thus too faint for spectral monitoring with the P60 Spectral Energy Distribution Machine (SEDM; Blagorodnova et al. 2018; Rigault et al. 2019) which is typically the workhorse for the BTS project. Instead, once the SN had rebrightened, we obtained four epochs of spectroscopy with the NOT using ALFOSC and three epochs with the Keck I telescope and the Low-Resolution Imaging Spectrometer (LRIS; Oke et al. 1995). A log of the spectral observations is provided in Table 2. The epochs of spectroscopy are also indicated by the downward-pointing arrows on top of the light curves in Fig. 2. The spectra obtained with ALFOSC and LRIS were reduced with the Pypeit (Prochaska et al. 2020) and LPipe (Perley 2019) software packages, respectively. The spectra were absolute calibrated against the band magnitudes using Gaussian Process (GP) interpolated magnitudes and then corrected for MW extinction. All spectral data will be made available via WISeREP222https://wiserep.weizmann.ac.il (Yaron & Gal-Yam 2012). We present the peak classification spectrum in Fig. 3 and the later sequence of spectra in Fig. 4.
@@ 3.3 Light curve
Based on the spectra, we have established that SN 2020aeuh is consistent with a SN Ia during the first peak, and is dominated by CSM interaction during the second peak (Sect. 3.2). We fit the , , and photometry with SN Ia template light curves using SNooPy444https://users.obs.carnegiescience.edu/cburns/SNooPyDocs/html/ (Burns et al. 2011) to determine the peak of the light curves in the different filters, and to see if the stretch of the light curve is consistent with the peak luminosity in accordance with the Phillips relation (Phillips 1993) for SNe Ia. The light curves (Fig. 6) are well fit by a model with a stretch of corresponding to an expected peak magnitude 0.15 mag fainter than what is observed, which is well within the range for normal Type Ia SNe. We also note that SN 2020aeuh was investigated in connection with the second data release of cosmological SNe Ia from ZTF (Rigault et al. 2025; Dimitriadis et al. 2025; Terwel et al. 2025). It yielded a good SALT2 fit with parameters555http://ztfcosmo.in2p3.fr/target/ZTF20acyroke mag and mag (, ). SNe Ia-CSM have peak magnitudes that are typically brighter than those of normal SNe Ia, in the range mag (Silverman et al. 2013a; Sharma et al. 2023). The -band light curve model of SN 2020aeuh peaks at mag, consistent with the subtype.
@@ 3.5 Host Analysis
We retrieved science-ready coadded images of the host galaxy of SN 2020aeuh from Sloan Digital Sky Survey DR 9 (SDSS; Ahn et al. 2012) and processed WISE images (Wright et al. 2010) from the unWISE archive (Lang 2014)666http://unwise.me. The unWISE images are based on the public WISE data and include images from NEOWISE-Reactivation mission R-7 (Mainzer et al. 2014; Meisner et al. 2017). We used the software package LAMBDAR (Lambda Adaptive Multi-Band Deblending Algorithm in R) (Wright et al. 2016) and tools presented by Schulze et al. (2021) to measure the brightness of the host galaxy. The host is well detected in all bands. We measure a brightness of mag in the SDSS band (corrected for MW reddening). The brightness in other filters is reported in Table 3.
@@ 4.2 Outside the optical regime
Mid-infrared (MIR) emission from the shock-heated dust (if present) has been interpreted as another proxy for CSM interaction. SN Ia-CSM 2018evt was detected in the MIR and proposed to show evidence for dust formation (Wang et al. 2024), and the compilation of Mo et al. (2025) showed that MIR detections by NEOWISE are not uncommon among the nearby SNe Ia-CSM. The image-subtraction pipeline, developed by De et al. (2025, in prep.), did not yield a NEOWISE detection, neither during the period when SN 2020aeuh was optically visible, nor in the subsequent years after its optical fading.
@@ 4.3 Missing H and He, nomenclature
For stripped-envelope SNe, the sequence of stripping from SN Ib to SN Ic has an equivalence among the CSM-interacting transients, where the subclasses of H-depleted SNe Ibn and He-depleted SNe Icn now have been proposed to be joined by the even more stripped Type Ien SN 2021yfj (Schulze et al. 2024).
@@ 4.5 Light-curve modelling
The fiducial value of gives . Furthermore, since the shock-crossing time is , the fiducial value of gives . Therefore, is estimated as M☉. We consider this option to be unrealistic given that the CSM is also devoid of both H and He.
@@
We used the Moriya (2023) prescription to fit the second peak of SN 2020aeuh. The parameters of the models were constrained to be consistent with the fact that here we are dealing with the explosion of a white dwarf (the model has previously mostly been used for Type IIn SNe). The parameters used for the theoretical light curve are as follows (see Moriya 2023, for definitions and details of the models): ejecta density structure with and , erg, M☉, and kinetic energy conversion efficiency (e.g. Fransson et al. 2014). Given that the model assumes the interaction to dominate the luminosity, we initialised this powering mechanism at 30 d (Sect. 4.1). This implies an evacuated region between the progenitor’s surface and the CSM. With an ejecta velocity of 12,000 km s-1, this gives cm. Similarly, assuming that the interaction ended 175–200 d after the explosion, based on the light-curve timescale, we have cm. We assumed cm2 g-1 for the CSM, as an opacity of H-free ionised matter.
@@
The best light-curve fit is shown in Fig. 8 (upper panel). Since the shape of the light curve is determined solely by the density profile, the choice of is not affecting the light curve morphology. The inferred values are M☉ for cm ( M☉ for cm). The inferred value of the CSM density slope is , regardless of the choice for . The lower panel of Fig. 8 shows the corresponding 1–3 confidence intervals.
@@ 4.6.1 Single degenerate
Regardless of the mechanism that leads to a thermonuclear explosion, the amount of points to a heavy nondegenerate companion. A star on the heavier end of intermediate-mass stars must have formed recently. The upper limit on the SFR estimated for the host (Sect. 3.5) seems to disfavour recent star formation. However the large distance to the host prevents more precise estimates of the SFR at the location of the explosion.
@@ 4.6.2 Double degenerate
A C/O-rich CSM formed during the merger of two WDs has been invoked in the past to explain 03fg-like SNe Ia (e.g. Dutta et al. 2022; Dimitriadis et al. 2022; Siebert et al. 2024). Those events, also coined super-Chandrasekhar, show absorption features of unburned carbon in their early spectra. Furthermore, some of the light curves show an early flux excess (Jiang et al. 2021; Dimitriadis et al. 2023; Srivastav et al. 2023). While a C/O CSM could explain those events (e.g. Noebauer et al. 2016), neither the distance nor the amount of CSM is similar to that required for SN 2020aeuh.
@@ 5 Summary and conclusions
During the early phase of its evolution, SN 2020aeuh appeared as a slightly overluminous SN Ia (Sect. 3.3). Its photometric properties during the first peak placed it marginally within the width-luminosity relationship of normal SNe Ia (Phillips 1993), near the outlying subtypes of 03fg-like, 91T-like, and Ia-CSM SNe (see Fig. 1 in Taubenberger 2017). While its later evolution revealed it as a SN Ia-CSM, its classification spectrum is best matched to 91T-like SNe. Furthermore, the H- and He-poor CSM is reminiscent of the postulated configuration for the progenitors of 03fg-like SNe, albeit with vastly different properties (Sect. 4.6.2). SN 2020aeuh exploded in the outskirts of its host galaxy, like some 03fg-like events (Khan et al. 2011). In contrast to other SNe Ia-CSM, the host of SN 2020aeuh is an early-type galaxy with a very low SFR (Sect. 3.5).
@@
SN 2020aeuh is a unique SN Ia-CSM that does not quite fit the picture of any suggested progenitor system for thermonuclear explosions (Sect. 4.6). The composition of the CSM could point toward a DD system, but the inferred masses call for a complete obliteration of a massive WD companion. Within the DD scenario, the case of multiple dynamical encounters can shed substantial amounts of material from the secondary WD (Sect. 4.6.2). However, both the we infer and the detached nature of the CSM are difficult to explain in this scenario. Perhaps a lower could be required in the case of asymmetric CSM. Additionally, the delay time between the stripping of the companion WD (either in the tidal-tail ejection or the fly-by encounter paradigms) and the explosion might account for the large radius of the CSM. Either way, a systematic study of the mechanisms that could produce some CSM in the DD scenario is still lacking. Regarding the SD (or CD) scenario, extensive stripping of the companion’s H and He layers is required, and detailed stellar-evolution scenarios leading up to such a configuration are hitherto missing (e.g. Liu et al. 2023). The remarkably low SFR of the host (Sect. 3.5) further weakens the nondegenerate cases. Altogether, the exotic appearance and properties of SN 2020aeuh pose crucial difficulties to the progenitor pathways of SNe Ia-CSM in particular and of SNe Ia in general.

### v2

@@
Identifying the progenitors of thermonuclear supernovae (Type Ia supernovae; SNe Ia) remains a key objective in contemporary astronomy. The rare subclass of SNe Ia-CSM that interacts with circumstellar material (CSM) allows for studies of the progenitor’s environment before explosion, and generally favours single-degenerate progenitor channels. The case of SN Ia-CSM PTF11kx clearly connected thermonuclear explosions with hydrogen-rich CSM-interacting events, and the more recent SN 2020eyj connected SNe Ia with helium-rich companion progenitors. Both of these objects displayed delayed CSM interaction which established their thermonuclear nature. Here we present a study of SN 2020aeuh, a Type Ia-CSM with delayed interaction. We analyse photometric and spectroscopic data that monitor the evolution of SN 2020aeuh and compare its properties with those of peculiar SNe Ia and core-collapse SNe. At early times, the evolution of SN 2020aeuh resembles a slightly overluminous SN Ia. Later, the interaction-dominated spectra develop the same pseudocontinuum seen in Type Ia-CSM PTF11kx and SN 2020eyj. However, the later-time spectra of SN 2020aeuh lack hydrogen and helium narrow lines. Instead, a few narrow lines could be attributed to carbon and oxygen. We fit the pseudobolometric light curve with a CSM-interaction model, yielding a CSM mass of M☉. We propose that SN 2020aeuh was a Type Ia supernova that eventually interacted with a dense medium which was deficient in both hydrogen and helium. Whereas previous SNe Ia-CSM constitute our best evidence for nondegenerate companion progenitors, the CSM around SN 2020aeuh is more difficult to understand. We include a hydrodynamical simulation for a double-degenerate dynamical collision to showcase that such a progenitor scenario could produce significant amounts of hydrogen-poor CSM, although likely not as much as the inferred CSM mass around SN 2020aeuh. It is clear that SN 2020aeuh challenges current models for stellar evolution leading up to a SN Ia explosion.
@@
Supernovae (SNe) interacting with circumstellar material (CSM) are classified based on their spectra, which are dominated by narrow ( km s-1) emission lines (e.g. Gal-Yam 2017). These emission lines are interpreted as signs of shock interaction between fast-moving SN ejecta and slow-moving CSM (see, e.g. Smith 2017, for a review). The most commonly observed subtypes of interacting SNe are Type IIn (Schlegel 1990; Nyholm et al. 2020) and Type Ibn (Hosseinzadeh et al. 2017), which have spectra dominated by hydrogen (H) and helium (He) emission lines, respectively. Recently, the classes of H/He-poor CSM-interacting stripped-envelope core-collapse supernovae (SNe Icn; Gal-Yam et al. 2022; Perley et al. 2022) and even a proposed SN Ien (Schulze et al. 2025), were also discovered.
@@ 2.1 Discovery, initial observations and classification
SN 2020aeuh (also known as ATLAS20bkbb, ZTF20acyroke) was first reported to the Transient Name Server (TNS111https://www.wis-tns.org/) in January 2021 by the Asteroid Terrestrial-impact Last Alert System (ATLAS) collaboration (Tonry et al. 2021), with their first detection from Christmas day 2020 December 25 (UTC dates are used throughout this paper; = 59208.6 MJD) at mag in the cyan band.
@@ 2.2 Photometry
Following the discovery, we obtained regular follow-up photometry with P48 until SN 2020aeuh faded below the detection limit, 180 days after the first detection. Initial follow-up observations were performed in the and bands, with later phases also covered in the band. This was accomplished with the ZTF camera (Dekany et al. 2020) on the P48. The -, -, and -band P48 data include points on 88, 86, and 23 separate nights, respectively.
The photometry from the P48 come from the ZTF forced-photometry service (Masci et al. 2019, 2023) and from the P60 using the pipeline described in Fremling et al. (2016). The photometric magnitudes of SN 2020aeuh are listed in Table 1. Additionally, we make use of the forced-photometry service from the ATLAS survey (Tonry et al. 2018; Smith et al. 2020; Shingles et al. 2021), which contains valuable photometry in the and bands.
@@ 2.3 Spectroscopy
The public SN Ia classification spectrum of SN 2020aeuh mentioned above was obtained 24 days after discovery (Weil et al. 2021). The SN never quite reached apparent magnitude 19, which is the cutoff for the Bright Transient Survey (BTS; Fremling et al. 2020; Perley et al. 2020). It was thus too faint for spectral monitoring with the P60 Spectral Energy Distribution Machine (SEDM; Blagorodnova et al. 2018; Rigault et al. 2019) which is typically the workhorse for the BTS project. Instead, once the SN had rebrightened, we obtained four epochs of spectroscopy with the NOT using ALFOSC and three epochs with the Keck I telescope and the Low-Resolution Imaging Spectrometer (LRIS; Oke et al. 1995). A log of the spectral observations is provided in Table 2. The epochs of spectroscopy are also indicated by the downward-pointing arrows on top of the light curves in Fig. 2. The spectra obtained with ALFOSC and LRIS were reduced with the Pypeit (Prochaska et al. 2020) and LPipe (Perley 2019) software packages, respectively. The spectra were absolute calibrated against the -band magnitudes using Gaussian Process (GP) interpolated magnitudes and then corrected for MW extinction. All spectral data are made available via WISeREP222https://www.wiserep.org/ (Yaron & Gal-Yam 2012). We present the peak classification spectrum in Fig. 3 and the later sequence of spectra in Fig. 4.
@@ 3.3 Light curve
Based on the spectra, we have established that SN 2020aeuh is consistent with a SN Ia during the first peak, and is dominated by CSM interaction during the second peak (Sect. 3.2). We fit the , , and photometry with SN Ia template light curves using SNooPy444https://users.obs.carnegiescience.edu/cburns/SNooPyDocs/html/ (Burns et al. 2011) to determine the peak of the light curves in the different filters, and to see if the stretch of the light curve is consistent with the peak luminosity in accordance with the Phillips relation (Phillips 1993) for SNe Ia. The light curves (Fig. 6) are well fit by a model with a stretch of corresponding to an expected peak magnitude 0.15 mag fainter than what is observed, which is well within the range for normal Type Ia SNe. We also note that SN 2020aeuh was investigated in connection with the second data release of cosmological SNe Ia from ZTF (Rigault et al. 2025; Dimitriadis et al. 2025; Terwel et al. 2025). It yielded a good SALT2 fit with parameters555http://ztfcosmo.in2p3.fr/target/ZTF20acyroke mag and mag (, ). SNe Ia-CSM have peak magnitudes that are typically brighter than those of normal SNe Ia, in the range mag (Silverman et al. 2013a; Sharma et al. 2023). The -band light-curve model of SN 2020aeuh peaks at mag, consistent with the subtype.
@@ 3.4 Pseudobolometric light curve
3.5 56Ni mass estimation
We can estimate the amount of 56Ni produced in SN 2020aeuh from its pseudobolometric light curve using the Arnett rule (Arnett 1982), which states that the peak bolometric luminosity equals the instantaneous total energy production rate from the radioactive decay of 56Ni, . Nadyozhin (1994) calculated to be
where d and d are the -folding timescales for the 56Ni and 56Co decays. Using and days, we obtain M☉. This is large for a normal SN Ia (e.g. Stritzinger et al. 2006; Bora et al. 2022), but is expected for a luminous SN 1991T-like object (Nugent et al. 1995; Hoeflich et al. 1996). There are also several caveats in this estimate, from the simple method of estimating the mass, the uncertainty of the bolometric light curve encapsulating all the flux, and any contribution from the circumstellar interaction to the first peak.
@@ 3.6 Host Analysis
We retrieved science-ready coadded images of the host galaxy of SN 2020aeuh from the Sloan Digital Sky Survey DR 9 (SDSS; Ahn et al. 2012) and processed WISE images (Wright et al. 2010) from the unWISE archive (Lang 2014)666http://unwise.me. The unWISE images are based on the public WISE data and include images from NEOWISE-Reactivation mission R-7 (Mainzer et al. 2014; Meisner et al. 2017). We used the software package LAMBDAR (Lambda Adaptive Multi-Band Deblending Algorithm in R) (Wright et al. 2016) and tools presented by Schulze et al. (2021) to measure the brightness of the host galaxy. The host is well detected in all bands. We measure a brightness of mag in the SDSS band (corrected for MW reddening). The brightness in other filters is reported in Table 3.
@@ 4.2 Outside the optical regime
Mid-infrared (MIR) emission from the shock-heated dust (if present) has been interpreted as another proxy for CSM interaction. SN Ia-CSM 2018evt was detected in the MIR and proposed to show evidence for dust formation (Wang et al. 2024), and the compilation of Mo et al. (2025) showed that MIR detections by NEOWISE are not uncommon among the nearby SNe Ia-CSM. The image-subtraction pipeline, developed by De et al. (in prep.), did not yield a NEOWISE detection, neither during the period when SN 2020aeuh was optically visible, nor in the subsequent years after its optical fading.
@@ 4.3 Missing H and He, nomenclature
For stripped-envelope SNe, the sequence of stripping from SN Ib to SN Ic has an equivalence among the CSM-interacting transients, where the subclasses of H-depleted SNe Ibn and He-depleted SNe Icn now have been proposed to be joined by the even more stripped Type Ien SN 2021yfj (Schulze et al. 2025).
@@ 4.5 Light-curve modelling
The fiducial value of (corresponding to the opacity of H-free ionised matter) gives . Furthermore, since the shock-crossing time is , the fiducial value of gives . Therefore, is estimated as M☉. We consider this option to be unrealistic given that the CSM is also devoid of both H and He.
@@
We used the Moriya (2023) prescription to fit the second peak of SN 2020aeuh, assuming the effect of diffusion to be negligible. The parameters of the models were constrained to be consistent with the fact that here we are dealing with the explosion of a white dwarf (the model has previously mostly been used for Type IIn SNe). The parameters used for the theoretical light curve are as follows (see Moriya 2023, for definitions and details of the models): ejecta density structure with and , erg, M☉, and kinetic energy conversion efficiency (e.g. Fransson et al. 2014). Given that the model assumes the interaction to dominate the luminosity, we initialised this powering mechanism at 30 d (Sect. 4.1). This implies an evacuated region between the progenitor’s surface and the CSM. With an ejecta velocity of 12,000 km s-1, this gives cm. Similarly, assuming that the interaction ended 175–200 d after the explosion, based on the light-curve timescale, we have cm.
@@
The best light-curve fit is shown in Fig. 8 (upper panel). Since the shape of the light curve is determined solely by the density profile, the choice of is not affecting the light curve morphology. The inferred values are M☉ for cm ( M☉ for cm). The inferred value of the CSM density slope is , regardless of the choice for . The total optical depth of CSM when assuming cm2 g-1 is , consistent with the assumption of neglecting the effect of diffusion. The lower panel of Fig. 8 shows the corresponding 1–3 confidence intervals.
@@ 4.6.1 Single degenerate
Regardless of the mechanism that leads to a thermonuclear explosion, the amount of points to a heavy nondegenerate companion. A star on the heavier end of intermediate-mass stars must have formed recently. The upper limit on the SFR estimated for the host (Sect. 3.6) seems to disfavour recent star formation. However the large distance to the host prevents more precise estimates of the SFR at the location of the explosion.
@@ 4.6.2 Double degenerate
A C/O-rich CSM formed during the merger of two WDs has been invoked in the past to explain SN 2003fg-like SNe Ia (e.g. Dutta et al. 2022; Dimitriadis et al. 2022; Siebert et al. 2024). Those events, also coined super-Chandrasekhar, show absorption features of unburned carbon in their early spectra. Furthermore, some of the light curves show an early flux excess (Jiang et al. 2021; Dimitriadis et al. 2023; Srivastav et al. 2023), similar only to their related SN 2002es-like SNe (Hoogendam et al. 2024). While a C/O CSM could explain SN 2003fg-like events (e.g. Noebauer et al. 2016), neither the distance nor the amount of CSM is similar to that required for SN 2020aeuh.
@@ 5 Summary and conclusions
During the early phase of its evolution, SN 2020aeuh appeared as a slightly overluminous SN Ia (Sect. 3.3). Its photometric properties during the first peak placed it marginally within the width-luminosity relationship of normal SNe Ia (Phillips 1993), near the outlying subtypes of SN 2003fg-like, SN 1991T-like, and Ia-CSM SNe (see Fig. 1 of Taubenberger 2017). While its later evolution revealed it as a SN Ia-CSM, its classification spectrum is best matched to SN 1991T-like SNe. Furthermore, the H- and He-poor CSM is reminiscent of the postulated configuration for the progenitors of SN 2003fg-like SNe, albeit with vastly different properties (Sect. 4.6.2). SN 2020aeuh exploded in the outskirts of its host galaxy, like some SN 2003fg-like events (Khan et al. 2011). In contrast to other SNe Ia-CSM, the host of SN 2020aeuh is an early-type galaxy with a very low SFR (Sect. 3.6).
@@
SN 2020aeuh is a unique SN Ia-CSM that does not quite fit the picture of any suggested progenitor system for thermonuclear explosions (Sect. 4.6). The composition of the CSM could point toward a DD system, but the inferred masses call for a complete obliteration of a massive WD companion. Within the DD scenario, the case of multiple dynamical encounters can shed substantial amounts of material from the secondary WD (Sect. 4.6.2). However, both the we infer and the detached nature of the CSM are difficult to explain in this scenario. Perhaps a lower could be required in the case of asymmetric CSM. Additionally, the delay time between the stripping of the companion WD (either in the tidal-tail ejection or the fly-by encounter paradigms) and the explosion might account for the large radius of the CSM. Either way, a systematic study of the mechanisms that could produce some CSM in the DD scenario is still lacking. Regarding the SD (or CD) scenario, extensive stripping of the companion’s H and He layers is required, and detailed stellar-evolution scenarios leading up to such a configuration are hitherto missing (e.g. Liu et al. 2023). The remarkably low SFR of the host (Sect. 3.6) further weakens the nondegenerate cases. Altogether, the exotic appearance and properties of SN 2020aeuh pose crucial difficulties to the progenitor pathways of SNe Ia-CSM in particular and of SNe Ia in general.
Data Availability
All spectra are available at the WISeREP, via https://www.wiserep.org/object/17853. The entire photometry table (Table 1) is available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via https://cdsarc.cds.unistra.fr/viz-bin/cat/J/A+A/704/A135.